# AuroraCap: Efficient, Performant Video Detailed Captioning and a New Benchmark

### v1

@@ Abstract
Video detailed captioning is a key task which aims to generate comprehensive and coherent textual descriptions of video content, benefiting both video understanding and generation. In this paper, we propose AuroraCap, a video captioner based on a large multimodal model. We follow the simplest architecture design without additional parameters for temporal modeling. To address the overhead caused by lengthy video sequences, we implement the token merging strategy, reducing the number of input visual tokens. Surprisingly, we found that this strategy results in little performance loss. AuroraCap shows superior performance on various video and image captioning benchmarks, for example, obtaining a CIDEr of 88.9 on Flickr30k, beating GPT-4V (55.3) and Gemini-1.5 Pro (82.2). However, existing video caption benchmarks only include simple descriptions, consisting of a few dozen words, which limits research in this field. Therefore, we develop VDC, a video detailed captioning benchmark with over one thousand carefully annotated structured captions. In addition, we propose a new LLM-assisted metric VDCscore for bettering evaluation, which adopts a divide-and-conquer strategy to transform long caption evaluation into multiple short question-answer pairs. With the help of human Elo ranking, our experiments show that this benchmark better correlates with human judgments of video detailed captioning quality. Code, docs, weight, benchmark and training data are all avaliable at website.
@@ 1 Introduction
The task of video detailed captioning involves generating comprehensive and coherent textual descriptions of video content, capturing not only the primary actions and objects but also intricate details, contextual nuances, and temporal dynamics. It has emerged as a critical area of research in computer vision and natural language processing, with significant implications for the fields of robotics Yang et al. (2023b); Du et al. (2024), ego-centric perception Grauman et al. (2022; 2023), embodied agents Zhao et al. (2023; 2024c; 2024d); Deng et al. (2023; 2024); Zhao et al. (2024b), and video editing Chai et al. (2023) and generation Bar-Tal et al. (2024). The challenges of video detailed captioning compared to past problems include the limited detailed caption data for training and evaluation, and also the lack of a good evaluation metric.
@@
Before the emergence of Large Language Models (LLMs), previous models could only generate very short and rough descriptions of videos Wang et al. (2022); Xu et al. (2023); Yan et al. (2022); Yang et al. (2023a). Although these models have been trained on web-scale video-text datasets (e.g., HowTo100M Miech et al. (2019) and VideoCC3M Nagrani et al. (2022)), their capabilities remain limited due to their smaller scale and the lack of world knowledge possessed by LLMs. Recently, researchers start to build more powerful large multimodal models (LMMs) upon pretrained LLMs (e.g., LLaVA Liu et al. (2024b), InstructBlip Dai et al. (2024b), InternVL Chen et al. (2023b)). These models typically use intermediate components (e.g., Q-Former Li et al. (2023a) or an MLP) to connect the pre-trained vision transformer (ViT) Dosovitskiy et al. (2020) and the LLM. Expanding from image-based LMMs to video-based LMMs is a natural progression, as videos can be viewed as sequences of frames. While most LMMs start with loading the pre-trained weights from image models and are further fine-tuned with video-text data, we find that LLaVA-like models can be easily adapted to a video one without any additional parameters but only with high-quality video-text instruction data for fine-tuning.
However, naive treatment of videos as a series of image frames can result in significant computational overhead and may cause a generalization of the length problem Wang et al. (2024b). To address these concerns, more specifically, to reduce the number of visual tokens, Video-LLaMA Li et al. (2023f) adapts the video Q-former, MovieChat Song et al. (2023) uses a memory bank, LLaMA-VID Li et al. (2023g) simply uses global pooling, and FastV Chen et al. (2024b) drops visual tokens by attention rank within LLM layers. In this paper, we present AuroraCap, adapting a simple yet efficient method called Token Merging Bolya et al. (2022), which is proved to be effective in image and video classification and editing tasks Li et al. (2023e). To be specific, we gradually combine similar tokens in a transformer layer using a bipartite soft matching algorithm to reduce the number of visual tokens. Following this pattern, our experiments show that we can use only 10% to 20% visual tokens compared to the original tokens generated by ViT with a marginal performance drop in various benchmarks. With this technique, it is easier to support higher-resolution and longer video sequence inputs for training and inference.
We present results on several widely used benchmarks, but find that existing video understanding benchmarks are either question-answer-based Song et al. (2023); Chen & Dolan (2011); Caba Heilbron et al. (2015); Xu et al. (2016); Xiao et al. (2021); Fu et al. (2024); Wu et al. (2024), which cannot demonstrate detailed descriptive abilities, or they provide descriptions that are too short, with only a few words Xu et al. (2016); Caba Heilbron et al. (2015) as shown in Table 1. Some large-scale datasets focus on specific domains such as ego-centric Grauman et al. (2023) or contain low-quality videos and annotations Bain et al. (2021). Therefore, we construct the VDC (Video Detailed Captions) benchmark, which contains over one thousand high-quality video-caption pairs spanning a wide range of categories, and the resulting captions encompass rich world knowledge, object attributes, camera movements, and crucially, detailed and precise temporal descriptions of events. We utilize GPT-4o as our recaption assistant with a hierarchical prompt design. To preserve as much information as possible from the videos and maintain temporal consistency, we implement a dense-frame extraction strategy. Using the dense frames as input, despite description of the whole video, we also generate high-quality captions from different aspects, including objective facts, backgrounds, camera angles and movements. Manual quality inspection is employed to ensure the quality of the video captions. While existing video-caption datasets Chen et al. (2024d); Ju et al. (2024); Li et al. (2024b) offer structured captions, VDC is the first benchmark focused on detailed video captioning, providing significantly longer and more detailed captions than others as shown in the Table 1 and Section B.
@@
MSVD Chen & Dolan (2011) 1,970 1,970 70,028 607,339 13,010 8.67
MSR-VTT Xu et al. (2016) 7,180 10,000 200,000 1,856,523 29,316 9.28
ActivityNet Krishna et al. (2017a) 20,000 100,000 100,000 1,340,000 15,564 13.40
S-MiT Monfort et al. (2021) Open 515,912 515,912 515,912 5,618,064 50,570 10.89
M-VAD Torabi et al. (2015) 92 48,986 55,905 519,933 18,269 9.30
MPII-MD Rohrbach et al. (2013) Movie 94 68,337 68,375 653,467 24,549 9.56
Youcook2 Zhou et al. (2018) Cooking 2,000 15,400 15,400 121,418 2,583 7.88
Charades Sigurdsson et al. (2016) 9,848 10,000 27,380 607,339 13,000 22.18
VATEX Wang et al. (2019) Human 41,300 41,300 413,000 4994,768 44,103 12.09
@@
We also introduce a novel evaluation metric specifically designed for detailed captioning task. Traditional metrics like METEOR Banerjee & Lavie (2005), CIDEr Vedantam et al. (2015), and BLEU Papineni et al. (2002), designed for machine translation or short captions, fail to evaluate detailed captions which contain rich semantic information. On the other side, an LLM-based evaluation metric is commonly used in visual question answering benchmarks Maaz et al. (2023); Wu et al. (2024), especially for those generated by VLMs Song et al. (2024; 2023). However, we observe that the LLM-based evaluation metric still struggles to differentiate the quality of detailed captions and tends to give lower scores. To address these challenges, we propose VDCscore, a novel captioning evaluation metric that leverages the reliability of large language models (LLMs) by evaluating short visual question-answer pairs. We first decompose the ground-truth caption into a set of concise question-answer pairs using the LLM, then generate corresponding responses from the predicted caption. Finally, the LLM is used to assess the accuracy of each response to provide an overall score. In particular, our paper makes the following contributions:
@@
  • In Section 2, we illustrate how we can reduce the number of tokens used for image or video input before injecting into LLM with marginal performance drop. Using these insights, we propose AuroraCap, which is shown to be the state-of-the-art video captioning model.
@@
  • In Section 3, we present VDC, the first benchmark for detailed video captioning, featuring over one thousand videos with significantly longer and more detailed captions. We comprehensively evaluate proprietary and open-source models using our proposed VDCscore metric.
@@ 2.1 Architecture
To effectively leverage the capabilities of both the pre-trained LLM and visual model, which is typically CLIP Radford et al. (2021) or DINO Oquab et al. (2023), LLaVA adapt a simple multi-layer perceptron (MLP) as projection layer to connect each patch tokens of image features into the word embedding space. The original LLaVA model is trained by a two-stage instruction-tuning procedure, which first pretraining projection layer for feature alignment and then finetuning end-to-end while freeze the visual encoder. Recent works like Prismatic VLMs Karamcheti et al. (2024) and Idefics2 Laurençon et al. (2024) further explore the design space of LLaVA architecture. We adapt some conclusion from those works for training the model.
@@
To increase the throughput of existing ViT models, Token Merging Bolya et al. (2022) is proposed to gradually combines similar tokens in a transformer to reduce the number of tokens passing through ViT models. Token Merging has been proven to be effective on image and video classification tasks even without the need for training. Token Merging is applied between the attention and MLP within each transformer block as:
@@
where $\bf s$ is the number of patches the token represents after token merging. We conduct frame-wise token merging in AuroraCap, the visualization of token merging can be found in Appendix C.
@@ 2.2 Training Recipe
Building upon the exploration in works like Prismatic VLMs Karamcheti et al. (2024), Idefics2 Laurençon et al. (2024), and InternVL Chen et al. (2023b), we further adopt a three-stage training strategy, which can be noted as Pretraining stage, Vision stage and Language stage. The training data used in each stage are shown in Table E3, Table E4 and Table E5. More training details including hyper-parameters selection and data preprocessing operation can be found in Appendix E.
@@
Similar to LLaVA, we first align visual features from the vision encoder ViT with the word embedding space of LLMs. To achieve this, we freeze the pretrained ViT and LLM, training solely the vision-language connector. Consistent with LLaVA-1.5 Li et al. (2024a), we employ a two-layer MLP as the projection layer and pretrain on 1.3M image-caption pairs. To optimize performance, we explore various combinations of the pre-trained ViT and LLM in Appendix D.
@@
Unlike LLaVA, we next unfreeze the pretrained ViT while freezing the LLM during vision stage and train with the public data among various computer vision tasks (e.g., captioning, object identification, classification, reasoning, VQA, and etc.) to get better generalization Huh et al. (2024). The motivation for doing this is that CLIP ViT usually performs poorly in aspects such as Orientation and Direction, Positional and Relational Context, Quantity and Count Tong et al. (2024b). However, since the most of the collected datasets lack high-quality and detailed corresponding language descriptions, the labels often consist of only a few words or a short phrase when converted to text. Therefore, unfreezing the language model at this stage is risky, as it may lead to a degradation in the performance of the language model.
@@ 2.3 Evaluation
In this section we evaluate AuroraCap on various tasks including image captioning, video captioning, and video question answering. Appendix F show detailed evaluation settings.
Table 2: Comparison of AuroraCap with LLM-based SoTA methods on image captioning benchmarks under zero-shot setting. The number in the upper right corner indicates the number of shots. Scores with bold indicate the best performance under zero-shot setting.
@@
C B@1 B@4 M R C B@1 B@4 M R C B@1 B@4 M R
LLaVA-1.5-7B Liu et al. (2023) 74.9 71.7 28.4 26.1 52.8 105.5 82.6 40.2 30.3 59.4 110.3 73.0 29.7 29.2 55.5
LLaVA-1.5-13B 79.4 73.6 30.2 26.6 53.9 109.2 84.2 42.4 30.6 60.3 115.6 74.6 31.5 29.4 56.5
LLaVA-1.6-7B Liu et al. (2024a) 68.4 69.6 26.6 23.2 50.3 88.4 73.8 34.8 25.9 54.6 99.9 67.7 28.4 25.5 52.4
LLaVA-1.6-13B 66.6 65.2 24.2 22.2 48.8 88.1 68.7 34.0 25.4 54.9 101.8 62.2 27.5 24.6 52.1
MiniCPM-V-3B Hu et al. (2023) 66.8 68.0 25.1 27.2 51.0 89.9 79.1 33.2 29.7 55.8 94.2 69.8 23.9 28.3 52.3
DeCap Li et al. (2023d) 56.7 -- 21.2 21.8 -- 42.7 -- -- -- -- 91.2 -- 24.7 25.0 --
Flamingo-80B Alayrac et al. (2022) 67.2 -- -- -- -- -- -- -- -- -- 84.3 -- -- -- --
Chameleon-34B Team (2024) 74.72 -- -- -- -- -- -- -- -- -- 120.22 -- -- -- --
@@ C B@1 B@4 M R C B@1 B@4 M R C B@1 B@4 M R
Image Captioning.
We evaluate AuroraCap using CIDEr (C), BELU-4 (B@4), BELU-1 (B@1), METEOR (M), and ROUGE-L (R) metric on Flickr Plummer et al. (2015), NoCaps Agrawal et al. (2019), and COCO-Cap Lin et al. (2014) benchmarks and compare it with LLM-based state-of-the-art methods as shown in Table 2. We show the performance of image based AuroraCap under zero-shot settings. Notice that these benchmarks all contain short captions consisting of a single sentence, so they only partially reflect the model’s performance. The performance mentioned in the rest of this paper refers to video-based AuroraCap.
Video Captioning.
Although the current video captioning benchmarks are only contains one-sentence captions, to compare with prior work, we similarly evaluate on these benchmarks. We evaluate AuroraCap using CIDer (C), BELU-4 (B@4), BELU-1 (B@1), METEOR (M), and ROUGE-L (R) metric on MSR-VTT Xu et al. (2016), VATEX Wang et al. (2019) and compare it with other methods as shown in Table 3.
Table 3: Comparison of AuroraCap with SoTA methods on existing video captioning benchmarks under zero-shot setting.
@@
C B@1 B@4 M R C B@1 B@4 M R
ZeroCap Tewel et al. (2022) 9.6 -- 2.9 16.3 35.4 -- -- -- -- --
DeCap Li et al. (2023d) 18.6 -- 14.7 20.4 -- 18.7 -- 13.1 15.3 --
PaLI-3 Chen et al. (2022) 21.3 -- -- -- -- -- -- -- -- --
Ma et al. Ma et al. (2024) 22.1 -- 3.5 17.3 28.7 23.9 -- 2.8 14.1 23.5
LLaVA-7B Liu et al. (2024b) 16.9 -- -- -- -- -- -- -- -- --
Video-LLaMA Zhang et al. (2023a) 2.3 -- 4.9 16.8 -- 3.8 -- 4.3 16.3 21.8
@@ C B@1 B@4 M R C B@1 B@4 M R
Table 4: Comparison of AuroraCap with SoTA methods on video question answering and classification benchmarks under zero-shot setting. The pretrained llm size is 7B for all the listed models.
@@
Just Ask Yang et al. (2021) -- -- -- -- -- -- 12.2
FrozenBiLM Yang et al. (2022) 24.7 -- 32.2 -- 16.8 -- 26.8
Video-LLaMA Zhang et al. (2023a) 12.4 1.1 51.6 2.5 29.6 1.8 --
VideoChat Li et al. (2023b) 26.5 2.2 56.3 2.8 45.0 2.5 --
Video-ChatGPT Maaz et al. (2023) 35.2 2.7 64.9 3.3 49.3 2.8 --
LLaMA-VID Li et al. (2023g) 47.4 3.3 69.7 3.7 57.7 3.2 --
Video-LLaVA Lin et al. (2023a) 45.3 3.3 70.7 3.9 59.2 3.5 --
FreeVA Wu (2024) 51.2 3.5 73.8 4.1 60.0 3.5 --
LLaVA-NeXT-Video Liu et al. (2024a) 53.5 3.2 -- -- -- -- --
MovieChat Song et al. (2023) 45.7 3.4 75.2 3.8 52.7 2.6 --
MovieChat+ Song et al. (2024) 48.1 3.4 76.5 3.9 53.9 2.7 --
@@
We evaluate AuroraCap on MSVD-QA Xu et al. (2017), ActivityNet-QA Yu et al. (2019), MSRVTT-QA Xu et al. (2017), and iVQA Yang et al. (2021) for video question answering tasks as shown in Table 4. Although AuroraCap is primarily a captioning model, it achieves competitive performance in some VQA datasets (ANet, iVQA). For others (MSVD, MSR-VTT) performance is more modest, but still not bad. In some failure cases observed in the model, we found that prompting the model to generate a comprehensive caption for the video input can lead to outputs that include the correct answer. This phenomenon may be attributed to a disruption in the model’s instruction-following capabilities during the visual-based training period. We regard this as a promising avenue for future research.

### v2

@@ Abstract
Video detailed captioning is a key task which aims to generate comprehensive and coherent textual descriptions of video content, benefiting both video understanding and generation. In this paper, we propose AuroraCap, a video captioner based on a large multimodal model. We follow the simplest architecture design without additional parameters for temporal modeling. To address the overhead caused by lengthy video sequences, we implement the token merging strategy, reducing the number of input visual tokens. Surprisingly, we found that this strategy results in little performance loss. AuroraCap shows superior performance on various video and image captioning benchmarks, for example, obtaining a CIDEr of 88.9 on Flickr30k, beating GPT-4V (55.3) and Gemini-1.5 Pro (82.2). However, existing video caption benchmarks only include simple descriptions, consisting of a few dozen words, which limits research in this field. Therefore, we develop VDC, a video detailed captioning benchmark with over one thousand carefully annotated structured captions. In addition, we propose a new LLM-assisted metric VDCscore for bettering evaluation, which adopts a divide-and-conquer strategy to transform long caption evaluation into multiple short question-answer pairs. With the help of human Elo ranking, our experiments show that this benchmark better correlates with human judgments of video detailed captioning quality.
@@ 1 Introduction
The task of video detailed captioning involves generating comprehensive and coherent textual descriptions of video content, capturing not only the primary actions and objects but also intricate details, contextual nuances, and temporal dynamics. It has emerged as a critical area of research in computer vision and natural language processing, with significant implications for the fields of robotics [148, 32], ego-centric perception [40, 41], embodied agents [163, 164, 165, 166, 29, 30], and video editing [13] and generation [7]. The challenges of video detailed captioning compared to past problems include the limited detailed caption data for training and evaluation, and also the lack of a good evaluation metric.
@@
Before the emergence of Large Language Models (LLMs), previous models could only generate very short and rough descriptions of videos [130, 139, 144, 147]. Although these models have been trained on web-scale video-text datasets (e.g., HowTo100M [93] and VideoCC3M [97]), their capabilities remain limited due to their smaller scale and the lack of world knowledge possessed by LLMs. Recently, researchers start to build more powerful large multimodal models (LMMs) upon pretrained LLMs (e.g., LLaVA [82], InstructBlip [27], InternVL [23]). These models typically use intermediate components (e.g., Q-Former [62] or an MLP) to connect the pre-trained vision transformer (ViT) [31] and the LLM. Expanding from image-based LMMs to video-based LMMs is a natural progression, as videos can be viewed as sequences of frames. While most LMMs start with loading the pre-trained weights from image models and are further fine-tuned with video-text data, we find that LLaVA-like models can be easily adapted to a video one without any additional parameters but only with high-quality video-text instruction data for fine-tuning.
@@
MSVD [14] 1,970 1,970 70,028 607,339 13,010 8.67
MSR-VTT [140] 7,180 10,000 200,000 1,856,523 29,316 9.28
ActivityNet [56] 20,000 100,000 100,000 1,340,000 15,564 13.40
S-MiT [96] Open 515,912 515,912 515,912 5,618,064 50,570 10.89
M-VAD [126] 92 48,986 55,905 519,933 18,269 9.30
MPII-MD [107] Movie 94 68,337 68,375 653,467 24,549 9.56
Youcook2 [168] Cooking 2,000 15,400 15,400 121,418 2,583 7.88
Charades [112] 9,848 10,000 27,380 607,339 13,000 22.18
VATEX [133] Human 41,300 41,300 413,000 4994,768 44,103 12.09
@@
However, naive treatment of videos as a series of image frames can result in significant computational overhead and may cause a generalization of the length problem [131]. To address these concerns, more specifically, to reduce the number of visual tokens, Video-LLaMA [72] adapts the video Q-former, MovieChat [114] uses a memory bank, LLaMA-VID [73] simply uses global pooling, and FastV [16] drops visual tokens by attention rank within LLM layers. In this paper, we present AuroraCap, adapting a simple yet efficient method called Token Merging [10], which is proved to be effective in image and video classification and editing tasks [71]. To be specific, we gradually combine similar tokens in a transformer layer using a bipartite soft matching algorithm to reduce the number of visual tokens. Following this pattern, our experiments show that we can use only 10% to 20% visual tokens compared to the original tokens generated by ViT with a marginal performance drop in various benchmarks. With this technique, it is easier to support higher-resolution and longer video sequence inputs for training and inference.
We present results on several widely used benchmarks, but find that existing video understanding benchmarks are either question-answer-based [114, 14, 11, 140, 137, 36, 134], which cannot demonstrate detailed descriptive abilities, or they provide descriptions that are too short, with only a few words [140, 11] as shown in Table 1. Some large-scale datasets focus on specific domains such as ego-centric [41] or contain low-quality videos and annotations [5]. Therefore, we construct the VDC (Video Detailed Captions) benchmark, which contains over one thousand high-quality video-caption pairs spanning a wide range of categories, and the resulting captions encompass rich world knowledge, object attributes, camera movements, and crucially, detailed and precise temporal descriptions of events. We utilize GPT-4o as our recaption assistant with a hierarchical prompt design. To preserve as much information as possible from the videos and maintain temporal consistency, we implement a dense-frame extraction strategy. Using the dense frames as input, despite description of the whole video, we also generate high-quality captions from different aspects, including objective facts, backgrounds, camera angles and movements. Manual quality inspection is employed to ensure the quality of the video captions. While existing video-caption datasets [19, 51, 61] offer structured captions, VDC is the first benchmark focused on detailed video captioning, providing significantly longer and more detailed captions than others as shown in the Table 1 and Section 8.
@@
We also introduce a novel evaluation metric specifically designed for detailed captioning task. Traditional metrics like METEOR [6], CIDEr [127], and BLEU [99], designed for machine translation or short captions, fail to evaluate detailed captions which contain rich semantic information. On the other side, an LLM-based evaluation metric is commonly used in visual question answering benchmarks [87, 134], especially for those generated by VLMs [115, 114]. However, we observe that the LLM-based evaluation metric still struggles to differentiate the quality of detailed captions and tends to give lower scores. To address these challenges, we propose VDCscore, a novel captioning evaluation metric that leverages the reliability of large language models (LLMs) by evaluating short visual question-answer pairs. We first decompose the ground-truth caption into a set of concise question-answer pairs using the LLM, then generate corresponding responses from the predicted caption. Finally, the LLM is used to assess the accuracy of each response to provide an overall score. In particular, our paper makes the following contributions:
@@
  • In Section 2, we illustrate how we can reduce the number of tokens used for image or video input before injecting into LLM with marginal performance drop. Using these insights, we propose AuroraCap, a video understanding base model with improved captioning performance, which also demonstrates advanced performance on existing benchmarks.
@@
  • In Section 3, we present VDC, the first benchmark for detailed video captioning, featuring over one thousand videos with significantly longer and more detailed captions. We comprehensively evaluate proprietary and open-source models using our proposed VDCscore metric. We also conduct extensive ablation studies based on AuroraCap.
@@ 2.1 Architecture
To effectively leverage the capabilities of both the pre-trained LLM and visual model, which is typically CLIP [103] or DINO [98], LLaVA adapt a simple multi-layer perceptron (MLP) as projection layer to connect each patch tokens of image features into the word embedding space. The original LLaVA model is trained by a two-stage instruction-tuning procedure, which first pretraining projection layer for feature alignment and then finetuning end-to-end while freeze the visual encoder. Recent works like Prismatic VLMs [52] and Idefics2 [58] further explore the design space of LLaVA architecture. We adapt some conclusion from those works for training the model.
@@
To increase the throughput of existing ViT models, Token Merging [10] is proposed to gradually combines similar tokens in a transformer to reduce the number of tokens passing through ViT models. Token Merging has been proven to be effective on image and video classification tasks even without the need for training. Token Merging is applied between the attention and MLP within each transformer block as:
@@
where $\bf s$ is the number of patches the token represents after token merging. We conduct frame-wise token merging in AuroraCap, the visualization of token merging can be found in Appendix 9.
@@ 2.2 Training Recipe
Building upon the exploration in works like Prismatic VLMs [52], Idefics2 [58], and InternVL [23], we further adopt a three-stage training strategy, which can be noted as Pretraining stage, Vision stage and Language stage. The training data used in each stage are shown in Table 13, Table 14 and Table 15. More training details including hyper-parameters selection and data preprocessing operation can be found in Appendix 11.
@@
Similar to LLaVA, we first align visual features from the vision encoder ViT with the word embedding space of LLMs. To achieve this, we freeze the pretrained ViT and LLM, training solely the vision-language connector. Consistent with LLaVA-1.5 [60], we employ a two-layer MLP as the projection layer and pretrain on 1.3M image-caption pairs. To optimize performance, we explore various combinations of the pre-trained ViT and LLM in Appendix 10.
@@
Unlike LLaVA, we next unfreeze the pretrained ViT while freezing the LLM during vision stage and train with the public data among various computer vision tasks (e.g., captioning, object identification, classification, reasoning, VQA, etc.) to get better generalization [46]. The motivation for doing this is that CLIP ViT usually performs poorly in aspects such as Orientation and Direction, Positional and Relational Context, Quantity and Count [125]. However, since the most of the collected datasets lack high-quality and detailed corresponding language descriptions, the labels often consist of only a few words or a short phrase when converted to text. Therefore, unfreezing the language model at this stage is risky, as it may lead to a degradation in the performance of the language model.
@@ 2.3 Evaluation
In this section we evaluate AuroraCap on various tasks including image captioning, video captioning, and video question answering. Appendix 12 show detailed evaluation settings.
Table 2: Comparison of AuroraCap with LLM-based SoTA methods on image-captioning benchmarks under a zero-shot setting, except $k$-shot where $k$ is marked as a superscript. The GPT-4V CIDEr results are from [119], the only reference we know of, but seem oddly low. * marks fine-tuned results. Scores in bold indicate the best performance under zero-shot setting.
@@
CI- BLEU BLEU Met- ROU- CI- BLEU BLEU Met- ROU- CI- BLEU BLEU Met- ROU-
DEr @1 @4 eor GE DEr @1 @4 eor GE DEr @1 @4 eor GE
InstructBLIP-7B [27] 78.2 77.1 30.9 24.8 53.4 118.2 88.2 47.2 30.3 62.0 141.3* 82.8* 41.7* 30.9* 61.0*
InstructBLIP-13B [27] 76.1 76.8 30.1 24.4 53.0 116.3 88.3 46.7 29.8 61.4 135.0* 82.2* 39.7* 29.8* 59.8*
LLaVA-1.5-7B [80] 74.9 71.7 28.4 26.1 52.8 105.5 82.6 40.2 30.3 59.4 110.3 73.0 29.7 29.2 55.5
LLaVA-1.5-13B [80] 79.4 73.6 30.2 26.6 53.9 109.2 84.2 42.4 30.6 60.3 115.6 74.6 31.5 29.4 56.5
LLaVA-1.6-7B [81] 68.4 69.6 26.6 23.2 50.3 88.4 73.8 34.8 25.9 54.6 99.9 67.7 28.4 25.5 52.4
LLaVA-1.6-13B [81] 66.6 65.2 24.2 22.2 48.8 88.1 68.7 34.0 25.4 54.9 101.8 62.2 27.5 24.6 52.1
MiniCPM-V-3B [43] 66.8 68.0 25.1 27.2 51.0 89.9 79.1 33.2 29.7 55.8 94.2 69.8 23.9 28.3 52.3
DeCap [67] 56.7 -- 21.2 21.8 -- 42.7 -- -- -- -- 91.2 -- 24.7 25.0 --
Flamingo-80B [2] 67.2 -- -- -- -- -- -- -- -- -- 84.3 -- -- -- --
Chameleon-34B [119] 74.72 -- -- -- -- -- -- -- -- -- 120.22 -- -- -- --
@@
Table 3: Comparison of AuroraCap with SoTA methods on existing video captioning benchmarks under zero-shot setting. AuroraCap substantially outperforms other recent methods.
@@
CI- BLEU BLEU Met- ROU- CI- BLEU BLEU Met- ROU-
DEr @1 @4 eor GE DEr @1 @4 eor GE
ZeroCap [121] 9.6 -- 2.9 16.3 35.4 -- -- -- -- --
DeCap [67] 18.6 -- 14.7 20.4 -- 18.7 -- 13.1 15.3 --
PaLI-3 [21] 21.3 -- -- -- -- -- -- -- -- --
Ma et al. [86] 22.1 -- 3.5 17.3 28.7 23.9 -- 2.8 14.1 23.5
LLaVA-7B [82] 16.9 -- -- -- -- -- -- -- -- --
Video-LLaMA [155] 2.3 -- 4.9 16.8 -- 3.8 -- 4.3 16.3 21.8
@@
Table 4: Comparison of AuroraCap with SoTA methods on video question answering and classification benchmarks under zero-shot setting. The pretrained LLM size is 7B for all listed models.
@@
Just Ask [145] -- -- -- -- -- -- 12.2
FrozenBiLM [146] 24.7 -- 32.2 -- 16.8 -- 26.8
Video-LLaMA [155] 12.4 1.1 51.6 2.5 29.6 1.8 --
VideoChat [64] 26.5 2.2 56.3 2.8 45.0 2.5 --
Video-ChatGPT [87] 35.2 2.7 64.9 3.3 49.3 2.8 --
LLaMA-VID [73] 47.4 3.3 69.7 3.7 57.7 3.2 --
Video-LLaVA [75] 45.3 3.3 70.7 3.9 59.2 3.5 --
FreeVA [136] 51.2 3.5 73.8 4.1 60.0 3.5 --
LLaVA-NeXT-Video [81] 53.5 3.2 -- -- -- -- --
MovieChat [114] 45.7 3.4 75.2 3.8 52.7 2.6 --
MovieChat+ [115] 48.1 3.4 76.5 3.9 53.9 2.7 --
@@
Image Captioning.
We evaluate AuroraCap using CIDEr (C), BLEU-4 (B@4), BLEU-1 (B@1), METEOR (M), and ROUGE-L (R) metric on Flickr [102], NoCaps [1], and COCO-Cap [78] benchmarks and compare it with LLM-based state-of-the-art methods as shown in Table 2. We show the performance of image based AuroraCap under zero-shot settings. Notice that these benchmarks all contain short captions consisting of a single sentence, so they only partially reflect the model’s performance. The performance mentioned in the rest of this paper refers to video-based AuroraCap.
Video Captioning.
Although the current video captioning benchmarks only contain one-sentence captions, to compare with prior work, we similarly evaluate on these benchmarks. We evaluate AuroraCap using CIDer (C), BLEU-4 (B@4), BLEU-1 (B@1), METEOR (M), and ROUGE-L (R) metric on MSR-VTT [140], VATEX [133] as shown in Table 3.
@@
We evaluate AuroraCap on MSVD-QA [138], ActivityNet-QA [152], MSRVTT-QA [138], and iVQA [145] for video question answering tasks as shown in Table 4. Although AuroraCap is primarily a captioning model, it achieves competitive performance in some VQA datasets (ANet, iVQA). For others (MSVD, MSR-VTT) performance is more modest, but still not bad. In some failure cases observed in the model, we found that prompting the model to generate a comprehensive caption for the video input can lead to outputs that include the correct answer. This phenomenon may be attributed to a disruption in the model’s instruction-following capabilities during the training.